OpenStudy (anonymous):

whats is the exact value of the expression square root of 9y^4+ 12square root of y^6-3square root y^6-the square root of y^4

OpenStudy (anonymous):

\[\sqrt{9y^4} + 12 \sqrt{y^6} - 3 \sqrt{y^6} - \sqrt{y^4}\]

OpenStudy (anonymous):

\[3y^2 + 12y^3 - 3y^3 - y^2\]

OpenStudy (anonymous):

so its 2y^2+9y^3?

OpenStudy (anonymous):

= \[9y^3 + 2y^2\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no prob. dont forget to presss best response button :P